# The development of a high granular crystal calorimeter prototype of VLAST

### v1

@@ I Introduction
Space-based gamma-ray astronomy offers unparalleled advantages for observing the universe. Unburdened by Earth’s atmosphere, these observatories enjoy broad bandwidth coverage, enabling detection of gamma rays across a vast energy spectrum, from MeV to TeV. This wide coverage is crucial for studying diverse phenomena, from nuclear processes in stars to the extreme environments around black holes and pulsars. The continuous monitoring capability of space-based telescopes provides excellent temporal resolution, allowing scientists to track the evolution of transient events like gamma-ray bursts and flares from active galactic nuclei. Moreover, the stable platform of space minimizes background noise, leading to high measurement precision and improved sensitivity for detecting faint sources. These combined advantages make space-based gamma-ray detection a crucial tool for investigating fundamental questions in astrophysics and cosmology, including the nature of dark matter, the origin of cosmic rays, and the mechanisms driving powerful astrophysical phenomena bib:1 ; bib:2 ; bib:3 .
Several successful gamma-ray missions have been successfully carried out around the world. The Energetic Gamma-ray Experiment Telescope (EGRET) bib:4 on the Compton Gamma-ray Observatory (CGRO) significantly advanced our understanding of high-energy gamma-ray sources. The Astro-rivelatore Gamma a Immagini Leggero (AGILE) bib:5 and the Fermi Large Area Telescope (Fermi-LAT) bib:6 ; bib:7 have further expanded our knowledge, provided detailed maps of the gamma-ray sky and revealed a wealth of new sources. The Dark Matter Particle Explorer (DAMPE) bib:8 ; bib:9 focuses on precise measurements of high-energy cosmic rays and gamma rays, contributing to the search for dark matter signatures.
Building upon these achievements, the Very Large Area gamma-ray Space Telescope (VLAST) bib:10 is proposed as the next-generation world-leading space-based gamma-ray observatory with significantly enhanced capabilities. With a sensitive area of approximately 10 m2, an order of magnitude larger than Fermi-LAT, VLAST will achieve unprecedented sensitivity across a wider energy range, enabling the detection of fainter sources and more detailed studies of known objects. Its superior energy resolution and precise particle track reconstruction capabilities will further enhance its scientific reach, allowing for more accurate measurements of energy spectra and search for potential the dark matter signatures.
@@
VLAST’s payload instrument comprises three key components working in concert, as shown in Fig. 1. The outermost at the top, the Anti-Coincidence Detector (ACD) bib:11 ; bib:12 , is constructed from plastic scintillators and serves as a veto system to discriminate against charged particle background. The ACD also measures the energy loss of charged particles, aiding in the identification of light nuclei, while providing a trigger signal for inner detectors.
Beneath the ACD lies the Silicon Tracker and low Energy gamma Detector (STED) bib:13 . Configured as a 2 × 2 array, each quadrant of the STED contains eight superlayers. Each superlayer integrates a thallium-doped cesium iodide (CsI(Tl)) detector and two double-sided silicon microstrip detectors bib:13 . The CsI(Tl) detectors play a dual role: they directly measure the energy of MeV gamma rays and act as a converter for high-energy gamma rays, producing electron-positron pairs. The innovative usage of CsI(Tl) as a converter, instead of denser materials like tungsten, improves the energy measurement of lower-energy gamma rays. The silicon microstrip detector provides high angular resolution gamma-ray tracking and incident particle impact characterization through charged particle trajectory reconstruction.
At the base of the instrument is the High-Energy Imaging Calorimeter (HEIC), the heart of VLAST’s high-energy measurements, which employs bismuth germanate (BGO) crystal as its primary sensitive material. The HEIC can accurately measure the characteristic profiles of the energy deposition of the secondary particles induced from electromagnetic shower and hadronic shower for efficiently distinguishing electrons (or gamma-ray) from hadrons. The calorimeter is designed to cover an extensive energy range from 0.1 GeV to 20 TeV for gamma photons and electrons, together with performance specifications of energy resolution better than 2 percent for 50 GeV photons, and the electron/proton separation capability should be better than 104, which is essential for separating gamma-ray signals from the background of cosmic-ray protons.
The HEIC serves as a critical component in VLAST detectors, primarily responsible for the precise energy measurements and particle identification, so its design is a critical aspect of VLAST’s development. Two potential design approaches are currently under consideration: (1) Following the DAMPE calorimeter design bib:14 , which utilizes orthogonally arranged elongated BGO crystal bars bib:15 . While the long bar design offers potential advantages including better energy resolution, the manufacturing of 1.2-meter-long, high-quality BGO crystals presents significant technical challenges, and (2) Implementing a high-granular design composed of cubic BGO crystals, which offers a more practical approach, and its finer granular scheme enables more accurate electromagnetic shower profile measurement of energy deposition.
@@ II Design and development of HEIC-Cube prototype
Several advanced calorimeter designs have been developed by the CALICE collaboration, showcasing different approaches to achieving high granularity. The silicon-tungsten (SiW) electromagnetic calorimeter utilizes silicon detectors interspersed with tungsten absorber plates bib:16 . Silicon detectors offer excellent spatial resolution, allowing for precise measurements of the shower development. The tungsten absorber provides the necessary material for the electromagnetic shower development. The scintillator-tungsten/copper (ScW) ECAL employs plastic scintillators and a tungsten-copper alloy absorber bib:17 . Scintillators offer fast response times and good light yield, while the tungsten-copper absorber provides a compact design. The analog hadron calorimeter (AHCAL) is based on plastic scintillators and iron absorbers bib:18 ; bib:19 ; bib:20 . Iron is a cost-effective absorber material for hadronic calorimetry. These sampling calorimeters, which alternate layers of active material (detector) and passive material (absorber), provide good imaging capabilities but often compromise with energy resolution due to the sampling fluctuations inherent in their design.
@@
In contrast, homogeneous calorimeters, which are constructed entirely of active material, offer the potential for both superior imaging and high-energy resolution. An example of this approach is the calorimeter design of the High Energy cosmic-Radiation Detection (HERD) experiment bib:21 ; bib:22 , which utilizes an array of small-sized heavy crystals, such as LYSO, as sensitive units. The absence of passive absorber material in homogeneous calorimeters minimizes sampling fluctuations, leading to superior imaging performance and improved energy resolution. The HERD calorimeter, with its high granularity and homogeneous design, is optimized for measuring the energy and direction of high-energy cosmic rays.
@@
The High Energy Imaging Calorimeter (HEIC) prototype described previously adopts a similar approach to HERD, using BGO crystals as the active material. However, a key difference lies in the readout scheme. While HERD employs fiber optic extraction to collect the scintillation light, the HEIC prototype utilizes fully embedded electronics. This approach offers several advantages, including improved light collection efficiency, which contributes to better energy resolution, and simplified integration of the readout electronics with the detector elements. This direct readout also reduces the potential for signal degradation associated with long optical fibers. The combination of high granularity, homogeneous crystal design, and fully embedded electronics makes the HEIC prototype a promising technology for future gamma-ray space observation.
@@
Our research group has developed a kind of high-granular BGO crystal calorimeter prototype designed for precise energy measurements of up to 50 GeV, and it is expected to exhibit a fine energy resolution for electrons and gamma-rays. To match the profile of the electromagnetic shower more accurately, the scalability in longitudinal depth is more significant than in horizontal direction. Therefore, the prototype consists of 10 longitudinal layers, each layer arranged in a 5 × 5 array of cubic BGO crystal with size of 3 cm × 3 cm × 3 cm. The calorimeter exhibits a total depth of approximately 27 radiation lengths, which is supposed sufficient for containing high-energy electromagnetic showers. The photoelectronic devices are avalanche photodiodes (APDs), offering advantages in quantum efficiency and compactness. The electronic boards for signal processing are embedded inside the calorimeter structure, minimizing signal path lengths and optimizing performance. This design aims for a balance between granularity, energy resolution, and compactness, making it suitable for various high-energy physics experiments.
@@ II.1 Sensitive unit
BGO crystals are widely applicable in high-energy physics experiments, including accelerator and space-based cosmic ray detection, due to several key properties. Their high density allows efficient interaction with high-energy particles, while non-hygroscopicity ensures stable performance in varying environmental conditions. BGO’s high light yield translates to better energy resolution, crucial for precise measurements. Furthermore, the crystals are relatively easy to produce and process into desired shapes, simplifying detector construction. The HEIC-Cube prototype employs a set of BGO scintillators with dimensions of 30 mm × 30 mm × 30 mm, and each surface of the crystal is encased in a 0.3 mm thick barium sulfate (BaSO4) reflective coating, as depicted in Fig. 2. This coating maximizes light collection by reflecting scintillation photons towards a designated 18 mm × 18 mm exit window on one surface of the cube, where a photodetector converts the light signal into an electrical pulse for subsequent analysis. This configuration optimizes light collection efficiency and relieves uneven light output to a certain extent, contributing to the detector’s overall performance.
When coupling with the emission wavelength of BGO scintillator, semiconductor photodetectors will have a better performance with higher quantum efficiency compared to photomultiplier tubes (PMTs) bib:23 . The improved efficiency translates to a stronger signal for a given amount of scintillation light. After comprehensive consideration of critical parameters including device dimension, gain characteristic, and linearity range, the HAMAMATSU S8664-0505 avalanche photodiode (APD) bib:24 was selected for this study. The S8664-0505 features a 5 mm × 5 mm active area and employs a P-on-N structure bib:25 , as shown in Fig. 2. The avalanche region P has an exceptionally thin profile of about 10 µm thick, which is crucial for minimizing unwanted signals from secondary particles. In this configuration, only the electrons in the P region generated through photoelectric conversion or ionization process can initiate the avalanche multiplication, amplifying the primary scintillation signal. In contrast, holes produced in the N- and N+ regions simply drift into the avalanche region without producing an avalanche. This unique structure effectively suppresses noise contributions from secondary particles generated by clustering effect directly within the N- and N+ regions of the APD. This targeted avalanche mechanism enhances the signal-to-noise ratio, improving the accuracy of energy measurements.
@@ II.2 Electronics
The readout scheme adopts waveform sampling mode, capturing the complete signal shape from each APD. This system consists of two primary components: the Pre-Amplifier Module (PAM), and the Analog-to-Digital Module (ADM) bib:26 , and they are interconnected via a FPGA Mezzanine Card (FMC) connector. This system performs pre-amplification and digitization of APD signals, as demonstrated in Fig. 4. Within the PAM, the JFETs are placed after each APD to suppress noise. A Pole-Zero Cancellation circuit (PZC) then shapes the Charge Sensitive Amplifier (CSA) output, which is subsequently split into high-gain and low-gain channels to extend the dynamic range. This dual-gain approach allows for accurate measurement of both small signals from minimum ionizing particles and larger signals from high-energy depositions.
The ADM houses a 12-bit, 32-channel Analog-to-Digital Converter (ADC) operating at 40 MSPS, digitizing the differential signals from all gain channels, and then they are transmitted to the FPGA for storage in an internal buffer. Each individual waveform consists of 512 sampling points (corresponding to 12.8 µs), ensuring that both the baseline and the entire waveform are completely captured within the sampling window. The FPGA in ADM also serves as instruction parsing and clock distribution. Additionally, a temperature sensor is placed next to the central APDs to monitor its operational temperature for calibration and performance analysis. This circuit also includes a Digital-to-Analog Converter (DAC) calibration mechanism. The linearity of the high-gain and low-gain channels was obtained respectively from the DAC calibration, which will inject a group of step signals with varying amplitudes to the CSA unit. The results show that the response coefficients of the high and low gain channels are 20.6 ADC/fC and 0.53 ADC/fC, with corresponding dynamic ranges are 150 fC and 7000 fC, respectively.
@@
The digitized signals are transmitted to a data acquisition board through optical fiber. The Data Concentrator Module (DCM) bib:27 primarily aggregates and buffers waveform data from multiple ADMs before uploading the consolidated data to the host computer via Ethernet for offline analysis and storage. Furthermore, the DCM incorporates essential functionalities including hit analysis, trigger generation, and command issuance. This hierarchical data acquisition system, from individual APDs to the central data collection, is designed for efficient and robust data handling in the laboratory environment. The inclusion of online data processing capabilities further optimizes data flow and reduces the volume of data transmitted to the back-end host.
@@ II.3 High dynamic range readout
The VLAST calorimeter is designed to detect primary cosmic rays with energies exceeding 20 TeV. Simulation results indicate that for electrons/photons of primary energy of 20 TeV, the deposited energy approach 7 TeV in a single crystal unit located in the shower center. The maximum energy in one crystal corresponds to approximately 2.4 × 105 MIPs, where MIP represents the energy deposition by a minimum ionization particle through a BGO crystal of 30 mm thickness, which is approximately 27.4 MeV based on simulations. To ensure accurate reconstruction of high-energy particle showers as well as reduce false triggering from electronic noise, the energy lower threshold is set to 0.1 MIPs for each detection unit. This low threshold is essential for capturing the full extent of the particle showers and distinguishing them from background noise. This condition necessitates the readout scheme of the calorimeter to achieve a dynamic range of 106.
@@
If a single readout device is placed under a BGO crystal for fluorescence collection, its dynamic range is limited by random noise and saturation effect, so it cannot satisfy the design specification of the large area calorimeter. To overcome these limitations, many international studies have developed photodetectors or readout electronics with multiple gain stages by providing progressive sensitivity levels to extend the dynamic range. For example, the calorimeter of Fermi-LAT adopts a pair of photodiodes with distinct sensitive areas bib:6 coupled to each end of individual scintillator. Furthermore, each of their readout electronics incorporate dual-gain channels to achieve additional dynamic range expansion. which ensures the effective energy range coverage from 20 MeV to about 300 GeV.
@@
Following the adoption route of semiconductor photodetectors, the HEIC employs a dual-APD configuration, where two APDs are positioned adjacent to each other in accordance with the scintillator geometry and coupled to the same fluorescence exit window of every BGO cubic units. These APDs are arranged directly facing the crystal’s light-emitting surface to ensure optimal fluorescence collection efficiency. A crucial element of the HEIC design is the incorporation of a light intensity attenuation filter. The entrance windows of one APD between the pair are covered with a fluorescence attenuation filter for extending the dynamic ranges for the high-energy end. This proportional attenuation allows the APD to remain within its linear operating range even when exposed to the intense light produced by high-energy particle interactions. The readout electronics used for both APDs are identical, and each APD leads out both an initial channel and an amplified channel on the back of the front-end board. This dual-APD configuration, combined with dual-gain channels in the readout electronics for each APD, provides the necessary dynamic range.
@@
To validate this large dynamic range readout scheme, an LED-based illumination system was established, as shown in Fig. 5. The readout system was enclosed in a darkroom environment to minimize stray light interference, with a LED positioned above the APD array. The LED served as a light source, simulating the scintillation light produced by the BGO crystal when interacting with cosmic rays. The pulse generator was used to excite LED, where the light intensity was precisely controlled through pulse voltage modulation bib:28 . The pulse duration was set to 300 ns which is consistent with the fluorescence decay time of BGO crystal. The signals from two APDs were independently processed by the readout electronics, with one APD covered by a filter with an attenuation coefficient about 1000 times placed on the incident window. Each APD signal is processed through two channels: a high-gain and a low-gain channel. This results in four readout channels per crystal: HH (High-gain from the unfiltered APD), HL (Low-gain from the unfiltered APD), LH (High-gain from the filtered APD), and LL (Low-gain from the filtered APD).
@@
During the testing process, the voltage of the pulse generator was continuously adjusted to modulate the LED light intensity, which allows for a systematic investigation of the readout system’s response across a wide range of simulated energy depositions. The gain ratios between adjacent readout channels were carefully measured and analyzed, as presented in Fig. 6. Linear fitting analysis revealed that the HH-HL and LH-LL ratios both exhibited values of approximately 36.5, which is determined by the design of the electronics. The HL-LH ratio was measured at about 31, a parameter introduced by the attenuation filter. Based on these ratios and the MIPs response characteristics of the BGO sensitive unit from Fig. 3 (showing a signal amplitude of 23.8 fC for a single MIP), the effective energy range of the four channel responses is detailed in Table 1. The combined dynamic range is evaluated as following: $\frac{4096-1200}{5×10}×36.36×31.39×36.76=2.4×10^{6}$, where the numerator on the left represents the effective range of HH channel, and the denominator represents the amplitude of the minimum effective signal, which equals to 5 times of the standard deviation of electronic noise. It will exceed 2.3 × 106 according to the relative gains between the readout channels, which proves to fulfill the design specifications for large-area calorimeter applications, demonstrating the effectiveness of the dual-APD, dual-gain readout scheme.

### v2

@@ I Introduction
Space-based gamma-ray astronomy offers unparalleled advantages for observing the universe. Unburdened by Earth’s atmosphere, these observatories enjoy broad bandwidth coverage, enabling detection of gamma rays across a vast energy spectrum, from MeV to TeV. This wide coverage is crucial for studying diverse phenomena, from nuclear processes in stars to the extreme environments around black holes and pulsars. The continuous monitoring capability of space-based telescopes provides excellent temporal resolution, allowing scientists to track the evolution of transient events like gamma-ray bursts and flares from active galactic nuclei. Moreover, the stable platform of space minimizes background noise, leading to high measurement precision and improved sensitivity for detecting faint sources. These combined advantages make space-based gamma-ray detection a crucial tool for investigating fundamental questions in astrophysics and cosmology, including the nature of dark matter, the origin of cosmic rays, and the mechanisms driving powerful astrophysical phenomena bib:1 ; bib:2 ; bib:3 ; bib:4 ; bib:5 .
Several successful gamma-ray missions have been successfully carried out around the world. The Energetic Gamma-ray Experiment Telescope (EGRET) bib:6 ; bib:7 on the Compton Gamma-ray Observatory (CGRO) significantly advanced our understanding of high-energy gamma-ray sources. The Astro-rivelatore Gamma a Immagini Leggero (AGILE) bib:8 and the Fermi Large Area Telescope (Fermi-LAT) bib:9 ; bib:10 ; bib:11 have further expanded our knowledge, provided detailed maps of the gamma-ray sky and revealed a wealth of new sources bib:12 ; bib:13 ; bib:14 ; bib:15 . The Dark Matter Particle Explorer (DAMPE) bib:16 ; bib:17 focuses on precise measurements of high-energy cosmic rays bib:18 ; bib:19 and gamma rays bib:5 ; bib:20 , contributing to the search for dark matter signatures.
Building upon these achievements, the Very Large Area gamma-ray Space Telescope (VLAST) bib:21 ; bib:22 ; bib:23 ; bib:24 ; bib:25 ; bib:31 is proposed as the next-generation world-leading space-based gamma-ray observatory with significantly enhanced capabilities. With a sensitive area of approximately 10 m2, an order of magnitude larger than Fermi-LAT, VLAST will achieve unprecedented sensitivity across a wider energy range, enabling the detection of fainter sources and more detailed studies of known objects. Its superior energy resolution and precise particle track reconstruction capabilities will further enhance its scientific reach, allowing for more accurate measurements of energy spectra and search for potential dark matter signatures.
@@
VLAST’s payload instrument comprises three key components working in concert, as shown in Fig. 1. The outermost detector at the top, the Anti-Coincidence Detector (ACD) bib:23 ; bib:24 , is constructed from plastic scintillators and serves as a veto system to discriminate against charged particle background. The ACD also measures the energy loss of charged particles, aiding in the identification of light nuclei, while providing a trigger signal for inner detectors.
Beneath the ACD lies the Silicon Tracker and low Energy gamma Detector (STED) bib:25 ; bib:26 . Configured as a 2 × 2 array, each quadrant of the STED contains eight superlayers. Each superlayer integrates a thallium-doped cesium iodide (CsI(Tl)) detector and two double-sided silicon microstrip detectors. The CsI(Tl) detectors play a dual role: they directly measure the energy of MeV gamma rays and act as a converter for high-energy gamma rays, producing electron-positron pairs. The innovative usage of CsI(Tl) as a converter, instead of denser materials like tungsten, improves the energy measurement of lower-energy gamma rays. The silicon microstrip detector provides high angular resolution gamma-ray tracking and incident particle impact characterization through charged particle trajectory reconstruction.
[…]
The HEIC serves as a critical component in VLAST detectors, primarily responsible for the precise energy measurements and particle identification, so its design is a critical aspect of VLAST’s development. Two potential design approaches are currently under consideration: (1) Following the DAMPE calorimeter design bib:27 ; bib:28 ; bib:29 ; bib:30 , which utilizes orthogonally arranged elongated BGO crystal bars bib:31 . While the long bar design offers potential advantages including better energy resolution, the manufacturing of 1.2-meter-long, high-quality BGO crystals presents significant technical challenges, and (2) Implementing a high-granular design composed of cubic BGO crystals, which offers a more practical approach, and its finer granular scheme enables more accurate electromagnetic shower profile measurement of energy deposition. The latter configuration shares many similarity with the CaloCube experiment bib:32 ; bib:33 ; bib:34 ; bib:35 , which was the first project consist of cubic crystals and has been validated to exhibit good performance.
@@ II Design and development of HEIC-Cube prototype
Several advanced calorimeter designs have been developed by the CALICE collaboration, showcasing different approaches to achieving high granularity. The silicon-tungsten (SiW) electromagnetic calorimeter utilizes silicon detectors interspersed with tungsten absorber plates bib:36 . Silicon detectors offer excellent spatial resolution, allowing for precise measurements of the shower development. The tungsten absorber provides the necessary material for the electromagnetic shower development. The scintillator-tungsten/copper (ScW) ECAL employs plastic scintillators and a tungsten-copper alloy absorber bib:37 ; bib:38 . Scintillators offer fast response times and good light yield, while the tungsten-copper absorber provides a compact design. The analog hadron calorimeter (AHCAL) is based on plastic scintillators and iron absorbers bib:39 ; bib:40 ; bib:41 ; bib:42 . Iron is a cost-effective absorber material for hadronic calorimetry. These sampling calorimeters, which alternate layers of active material (detector) and passive material (absorber), provide good imaging capabilities but often compromise with energy resolution due to the sampling fluctuations inherent in their design.
@@
In contrast, homogeneous calorimeters, which are constructed entirely of active material, offer the potential for both superior imaging and better energy resolution. An example of this approach is the calorimeter design of the High Energy cosmic-Radiation Detection (HERD) experiment bib:43 ; bib:44 ; bib:45 , which utilizes an array of small-sized heavy crystals, LYSO, as sensitive units. The absence of passive absorber material in homogeneous calorimeters minimizes sampling fluctuations, leading to superior imaging performance and improved energy resolution. The HERD calorimeter, with its high granularity and homogeneous design, is optimized for measuring the energy and direction of high-energy cosmic rays.
@@
The High Energy Imaging Calorimeter (HEIC) prototype described previously adopts a similar approach to HERD, using BGO crystals as the active material. However, a key difference lies in the readout scheme. HERD employs two independent systems bib:46 to collect the scintillation light: the first one uses wavelength shifting fibers bib:47 to deliver the light to Intensified scientific CMOS (IsCMOS) cameras, whereas the second one makes use of photo-diode sensors bib:48 ; bib:49 ; bib:50 . Avalanche photodiodes (APDs) are selected for the HEIC prototype, and their associated electronic board are embedded into the calorimeter. The configuration shares a series of similarities with the HERD photodiode readout scheme, which has been validated in its previous experiment bib:51 ; bib:52 . This approach offers several advantages, including simplified integration of the readout electronics with the detector elements and reduced system complexity. This direct readout also reduces the potential for signal degradation associated with long optical fibers, and it improves light collection efficiency, which contributes to better energy resolution. The combination of high granularity, homogeneous crystal design, and fully embedded electronics makes the HEIC prototype a promising technology for future gamma-ray space observation.
@@
We have developed a kind of high-granular BGO crystal calorimeter prototype designed for precise energy measurements of up to 50 GeV, and it is expected to exhibit a fine energy resolution for electrons and gamma-rays. To match the profile of the electromagnetic shower more accurately, the scalability in longitudinal depth is more significant than in horizontal direction. Therefore, the prototype consists of 10 longitudinal layers, each layer arranged in a 5 × 5 array of cubic BGO crystal with size of 3 cm × 3 cm × 3 cm. The calorimeter exhibits a total depth of approximately 27 radiation lengths, which is supposed sufficient for containing high-energy electromagnetic showers. The photoelectronic devices we chose are APDs, offering advantages in quantum efficiency and compactness when compared with photomultiplier tube (PMT) bib:53 . The electronic boards for signal processing are embedded inside the calorimeter structure, minimizing signal path lengths and optimizing performance. This design aims for a balance between granularity, energy resolution, and compactness, making it suitable for various high-energy physics experiments.
@@ II.1 Sensitive unit
BGO crystals are widely applicable in high-energy physics experiments, including accelerator and space-based cosmic ray detection bib:54 , due to several key properties. Their high density allows efficient interaction with high-energy particles, while non-hygroscopicity ensures stable performance in varying environmental conditions. BGO’s high light yield translates to better energy resolution, crucial for precise measurements. A group of inherent properties is listed in table 1. Furthermore, the crystals are relatively easy to produce and process into desired shapes, simplifying detector construction. The HEIC-Cube prototype employs a set of BGO scintillators with dimensions of 30 mm × 30 mm × 30 mm, and each surface of the crystal is encased in a 0.3 mm thick barium sulfate (BaSO4) reflective coating, as depicted in Fig. 2. This coating maximizes light collection by reflecting scintillation photons towards a designated 18 mm × 18 mm exit window on one surface of the cube, where a photodetector converts the light signal into an electrical pulse for subsequent analysis. This configuration optimizes light collection efficiency and relieves uneven light output to a certain extent, contributing to the detector’s overall performance.
Table 1: The properties of BGO [(Bi2O3)2(GeO2)3]
Quantity Value Units
Density 7.130 g cm-3
Radiation length 1.118 cm
Molière radius 2.259 cm
Minimum ionization 8.918 MeV cm-1
Nuclear interaction length 22.32 cm
Fluorescence emission wavelength peak 480 nm
Fluorescence decay time 300 ns
When coupling with the emission wavelength of BGO scintillator, semiconductor photodetectors will have a better performance with higher quantum efficiency compared to PMTs bib:53 . The improved efficiency translates to a stronger signal for a given amount of scintillation light. After comprehensive consideration of critical parameters including device dimension, gain characteristic, and linearity range, the HAMAMATSU S8664 - 0505 avalanche photodiode (APD) bib:55 illustrated in Fig. 2 was selected for this study. The S8664-0505 features a 5 mm × 5 mm active area and employs a P-on-N structure bib:56 , as shown in Fig. 2. The avalanche region P has an exceptionally thin profile of about 10 µm thick, which is crucial for minimizing unwanted signals from secondary particles. In this configuration, only the electrons in the P region generated through photoelectric conversion or ionization process can initiate the avalanche multiplication, amplifying the primary scintillation signal. In contrast, holes produced in the N- and N+ regions simply drift into the avalanche region without producing an avalanche. This unique structure effectively suppresses noise contributions from secondary particles generated by clustering effect directly within the N- and N+ regions of the APD. This targeted avalanche mechanism enhances the signal-to-noise ratio, improving the accuracy of energy measurements.
@@ II.2 Electronics
The readout scheme adopts waveform sampling mode, capturing the complete signal shape from each APD. This system consists of two primary components: the Pre-Amplifier Module (PAM), and the Analog-to-Digital Module (ADM) bib:57 ; bib:58 , and they are interconnected via a FPGA Mezzanine Card (FMC) connector. This system performs pre-amplification and digitization of APD signals, as demonstrated in Fig. 4. Within the PAM, the JFETs are placed after each APD to suppress noise. A Pole-Zero Cancellation circuit (PZC) then shapes the Charge Sensitive Amplifier (CSA) output, which is subsequently split into high-gain and low-gain channels to extend the dynamic range. This dual-gain approach allows for accurate measurement of both small signals from minimum ionizing particles and larger signals from high-energy depositions.
The ADM houses a 12-bit, 32-channel Analog-to-Digital Converter (ADC) operating at 40 MSPS, digitizing the differential signals from all gain channels, and then they are transmitted to the FPGA for storage in an internal buffer. Each individual waveform consists of 512 sampling points (corresponding to 12.8 µs), ensuring that both the baseline and the entire waveform are completely captured within the sampling window. The FPGA in ADM also serves as instruction parsing and clock distribution. Additionally, a temperature sensor is placed next to the central APDs to monitor its operational temperature for calibration and performance analysis. This circuit also includes a Digital-to-Analog Converter (DAC) calibration mechanism. The linearity of the high-gain and low-gain channels was obtained respectively from the DAC calibration, which will inject a group of step signals with varying amplitudes to the CSA unit. The results present that the response coefficients of the high and low gain channels are 20.6 ADC/fC and 0.53 ADC/fC, with corresponding dynamic ranges are 150 fC and 7000 fC, respectively.
@@
The digitized signals are transmitted to a data acquisition board through optical fiber. The Data Concentrator Module (DCM) bib:59 primarily aggregates and buffers waveform data from multiple ADMs before uploading the consolidated data to the host computer via Ethernet for offline analysis and storage. Furthermore, the DCM incorporates essential functionalities including hit analysis, trigger generation, and command issuance. This hierarchical data acquisition system, from individual APDs to the central data collection, is designed for efficient and robust data handling in the laboratory environment. The inclusion of online data processing capabilities further optimizes data flow and reduces the volume of data transmitted to the back-end host.
@@ II.3 High dynamic range readout
The VLAST calorimeter is designed to detect primary cosmic rays with energies exceeding 20 TeV. Simulation results indicate that for electrons/photons of primary energy of 20 TeV, the deposited energy approaches 7 TeV in a single crystal unit located in the shower center. The maximum energy in one crystal corresponds to approximately 2.4 × 105 MIPs, where MIP represents the energy deposition by a minimum ionization particle through a BGO crystal of 30 mm thickness, which is approximately 27.4 MeV based on simulations. To ensure accurate reconstruction of high-energy particle showers as well as reduce false triggering from electronic noise, the energy lower threshold is set to 0.1 MIPs for each detection unit. This low threshold is essential for capturing the full extent of the particle showers and distinguishing them from background noise. This condition necessitates the readout scheme of the calorimeter to achieve a dynamic range of 106.
@@
If a single readout device is placed under a BGO crystal for fluorescence collection, its dynamic range is limited by random noise and saturation effect, so it cannot satisfy the design specification of the large area calorimeter. To overcome these limitations, many international studies have developed photodetectors or readout electronics with multiple gain stages by providing progressive sensitivity levels to extend the dynamic range. For example, the calorimeter of Fermi-LAT adopts a pair of photodiodes with distinct sensitive areas bib:9 ; bib:10 coupled to each end of individual scintillator. Furthermore, each of their readout electronics incorporate dual-gain channels to achieve additional dynamic range expansion, which ensures the effective energy range coverage from 20 MeV to about 300 GeV.
@@
Following the adoption route of semiconductor photodetectors, the HEIC employs a dual-APD configuration, where two APDs are positioned adjacent to each other in accordance with the scintillator geometry and coupled to the same fluorescence exit window of every BGO cubic units. These APDs are arranged directly facing the crystal’s light-emitting surface to ensure optimal fluorescence collection efficiency. A crucial element of the HEIC design is the incorporation of a light intensity attenuation filter. The entrance windows of one APD between the pair are covered with a fluorescence attenuation filter for extending the dynamic ranges for the high-energy end. The attenuation filter, an exposure film, is made of a plastic sheet with a thickness of 0.17 mm. It is designed to tune the amplitudes of signals by proportional attenuating the intensity of scintillation light injected into the APDs, which remains it within the linear operating range of the electronic amplifiers. The readout electronics used for both APDs are identical, and each APD leads out both an initial channel and an amplified channel on the back of the front-end board. This dual-APD configuration, combined with dual-gain channels in the readout electronics for each APD, provides the necessary dynamic range.
@@
To validate this large dynamic range readout scheme, an LED-based illumination system was established, as depicted in Fig. 5. The readout system was enclosed in a darkroom environment to minimize stray light interference, with a LED positioned above the APD array. The LED served as a light source, simulating the scintillation light produced by the BGO crystal when interacting with cosmic rays. The pulse generator was used to excite LED, where the light intensity was precisely controlled through pulse voltage modulation bib:60 . The pulse duration was set to 300 ns which is consistent with the fluorescence decay time of BGO crystal. The signals from two APDs were independently processed by the readout electronics, with one APD covered by a filter with an attenuation coefficient about 1000 times placed on the incident window. Each APD signal is processed through two channels: a high-gain and a low-gain channel. This results in four readout channels per crystal: HH (High-gain from the unfiltered APD), HL (Low-gain from the unfiltered APD), LH (High-gain from the filtered APD), and LL (Low-gain from the filtered APD).
@@
During the testing process, the voltage of the pulse generator was continuously adjusted to modulate the LED light intensity, which allows for a systematic investigation of the readout system’s response across a wide range of simulated energy depositions. The gain ratios between adjacent readout channels were carefully measured and analyzed, as presented in Fig. 6. Linear fitting analysis revealed that the HH-HL and LH-LL ratios both exhibited values of approximately 36.5, which is determined by the design of the electronics. Figure 6 and 6 depict their amplitude correlations between high-gain and low-gain channels respectively. The HL-LH ratio was measured at about 31, a parameter introduced by the attenuation filter, illustrated in Fig. 6. Based on these ratios and the MIPs response characteristics of the BGO sensitive unit from Fig. 3 (showing a signal amplitude of 23.8 fC for a single MIP), the effective energy range of the four channel responses is detailed in Table 2. The combined dynamic range will exceed 2.4 × 106 according to the relative gains between the readout channels, which proves to fulfill the design specifications for large-area calorimeter applications, demonstrating the effectiveness of the dual-APD, dual-gain readout scheme.